Session: May 2004 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Xinyong Huang at Thu May 27 02:52:48 2004) Success :
we lost the scan No021 because of the antenna control PC problem. We stop it at UT:148.01:43, because of the sources are out of antenna limitation.

Onsala (Michael Lindqvist at Thu May 27 03:48:38 2004) Success :
No known problems

Jodrell_Bank (Alastair Gunn at Thu May 27 05:37:58 2004) Success :
The Lovell Telescope was used. No known problems.

Torun (K. Borkowski at Thu May 27 06:07:34 2004) Minor failures :
Operators noted that several scans were recorded with Tcal on. No other known problems.

Medicina (G.Maccaferri at Thu May 27 07:26:06 2004) Success :

Effelsberg (A. Kraus at Thu May 27 08:23:02 2004) Success :

Noto (P. Cassaro at Thu May 27 13:32:33 2004) Success :

Westerbork (Tony Foley at Fri May 28 08:54:37 2004) Minor failures : HA-E
Wb used in phased array mode. Some scans out of HA range at start replaced with local calibration. Otherwise looked fine